# FaceTalk: Audio-Driven Motion Diffusion for Neural Parametric Head Models

### 13 Dec 2023 v1

@@ Abstract
We introduce FaceTalk, a novel generative approach designed for synthesizing high-fidelity 3D motion sequences of talking human heads from input audio signal. To capture the expressive, detailed nature of human heads, including hair, ears, and finer-scale eye movements, we propose to couple speech signal with the latent space of neural parametric head models to create high-fidelity, temporally coherent motion sequences. We propose a new latent diffusion model for this task, operating in the expression space of neural parametric head models, to synthesize audio-driven realistic head sequences. In the absence of a dataset with corresponding NPHM expressions to audio, we optimize for these correspondences to produce a dataset of temporally-optimized NPHM expressions fit to audio-video recordings of people talking. To the best of our knowledge, this is the first work to propose a generative approach for realistic and high-quality motion synthesis of volumetric human heads, representing a significant advancement in the field of audio-driven 3D animation. Notably, our approach stands out in its ability to generate plausible motion sequences that can produce high-fidelity head animation coupled with the NPHM shape space. Our experimental results substantiate the effectiveness of FaceTalk, consistently achieving superior and visually natural motion, encompassing diverse facial expressions and styles, outperforming existing methods by 75% in perceptual user study evaluation [^1] Project Page: https://shivangi-aneja.github.io/projects/facetalk.
Figure 1: Given input speech signal, we propose a diffusion-based approach to synthesize high-quality and temporally consistent 3D motion sequences of high-fidelity human heads as neural parametric head models. Our method can generate a diverse set of expressions (including wrinkles and eye blinks) and the generated mouth motion is temporally synchronized with the given audio signal.
@@ 1 Introduction
Modeling 3D animation of humans has a wide range of applications in the realm of digital media, including animated movies, computer games, and virtual agents. In recent years, there have been numerous works proposing generative approaches for motion synthesis of human bodies, enabling the animation of human skeletons conditioned on various signals such as action [28, 51, 3], language [1, 52, 70, 82, 40, 4] and music [73, 2]. While human faces are critical to synthesis of humans, generative synthesis of 3D faces in motion has focused on 3D morphable models (3DMMs) leveraging linear blendshapes [7, 44] to represent head motion and expression. Such models characterize a disentangled space of head shape and motion, but lack the capacity to comprehensively represent the complexity and fine-grained details of human face geometry in motion (e.g., hair, skin furrowing during motion, etc.).
Thus, we propose to represent animated head sequences with a volumetric 3D head representation, leveraging the expressive representation space of neural parametric head models (NPHMs) [27]. NPHMs offer a flexible representation capable of handling complex and irregular facial expressions (e.g., blinking, skin creasing), along with a high-fidelity shape space including the head, hair, and ears, making them a much more suitable choice for face animation. We address the challenging task of creating an audio-conditional generative animation model for this volumetric representation.
@@ 3 Preliminaries
In contrast to traditional 3D morphable models, which remain limited in expressive detail, we employ the Neural Parametric Head Model (NPHM) representation [27]. Similar to traditional 3DMMs, NPHM disentangles a human face into an identity and expression space; however, the shape and expression space are volumetric, enabling expressive capture of details such as fine-scale eye movements, hair, ears, etc. NPHM uses two auto-decoder style neural networks (a) Identity Network ${}\big{\{}\mathcal{I}\big{\}}$ to represent the overall facial shape and (b) Expression Network ${}\mathcal{E}\big{\}}$ to represent the facial movements such as jaw pose, wrinkles, eyeblinks, etc., jointly denoted as $\mathcal{F}=\big{\{}\mathcal{I},\mathcal{E}\big{\}}$. The identity and expressions latent codes $\big{\{}\boldsymbol{{\theta}_{id}},\boldsymbol{{\theta}_{exp}}\big{\}}$ define the facial shape and expression respectively. Unlike 3DMMs which predict a fixed-topology mesh that can not model fine surface details, NPHM can handle different hairstyles, wrinkles and complex facial expressions. NPHM represents identities with a signed distance field (SDF) in canonical space and models the expressions as deformations onto the facial region. Mathematically, it can be defined as:
@@
Diffusion models are a class of generative models that consists of a forward and reverse process. The forward process converts the original structured data distribution into Gaussian noise, modeled following a fixed Markov chain as:
@@ 4 Method
FaceTalk performs high-fidelity and temporally consistent generative synthesis of motion sequences of heads, conditioned on audio signal. In order to characterize complex face motions and fine-scale movements, we synthesize realistic heads in the latent expression space of a parametric model for volumetric head representations, i.e., neural parametric head models (NPHMs) [27]. We, thus, develop a speech-conditioned latent diffusion model to synthesize temporally coherent head expression sequences that are coupled with the NPHM shape space to produce complex, realistic head animation of different identities. An overview of our approach is illustrated in Fig. 2.
@@
We train our expression decoder network on optimized NPHM expression sequences $\big{\{}\boldsymbol{\theta}_{exp}\big{\}}^{1:N}$ (obtained in Section 5), where $N$ refers to number of frames in a sequence. We train a diffusion-based stacked multi-layer transformer [74] decoder network to synthesize facial expressions in the latent space of the NPHM model. During training, following forward diffusion (Eq. 3) we add noise for a randomly sampled diffusion timestamp $t\sim\mathtt{Uniform}(0,T)$ to create noisy expression codes $\big{\{}\boldsymbol{\theta}_{exp}\big{\}}^{1:N}_{t}$. These noisy expression codes are then projected to the latent space of our model via a linear layer, followed by a stack of transformer decoder blocks, and then projected back to the original NPHM space using another linear layer. To embed diffusion timestamp in the latent space of our model, we apply sinusoidal embedding and process it through three-layer MLP. Next, to fuse diffusion timestamp into the model, we use a one-layer FiLM (feature-wise linear modulation) network [50] between the multi-head self-attention, multi-head cross-attention and feedforward layers of the transformer decoder block, which is critical for the model to produce high-quality expressions as we show in the results. We leverage a look-ahead binary target mask $\mathcal{T}\in\mathbb{R}^{N\times N}$ in the multi-head self-attention layers to prevent model from peeking into the future expression codes. Mathematically, it can be written as:
@@
where $\mathcal{T}_{ij}$ refers to the $(i,j)^{th}$ element of the matrix and $1\leq i,j\leq N$. The audio features $\boldsymbol{A}^{1:N}$ are fused into the network via the multi-head cross-attention blocks. We leverage the expression-audio alignment mask $\mathcal{M}$ to fuse the audio features into the network to ensure that mouth poses learned by the expression codes are consistent with the speech signal. The binary mask $\mathcal{M}\in\mathbb{R}^{N\times N}$ is Kronecker delta function $\delta_{ij}$ such that the audio features for $i^{th}$ timestamp attend to expression features at the $j^{th}$ timestamp if and only if $i=j$. Mathematically,
@@
Unlike other domains (like text-to-motion) where the synthesized motion can vary dramatically, speech-driven 3D facial animation is more constrained, requiring precise mouth alignment with the audio signal, making it prone to overfitting. To alleviate this, we propose a novel augmentation strategy capable of producing diverse expressions for a given audio signal. Our key insight is to augment the dataset to generate different expression codes for the same speech signal by randomly amplifying and suppressing its magnitude. Specifically, we randomly sample modulation factor $r$ as:
$r\sim\mathtt{Uniform}(0.5,3),$ (8)
and scale the expression codes as $\big{\{}r\times\boldsymbol{\theta}_{exp}\big{\}}^{1:N}$. We show in the results this augmentation strategy helps the model to synthesize diverse expressions.
@@
FaceTalk uses a diffusion-based framework to learn to synthesize NPHM expression sequences of $N$ frames $\big{\{}\boldsymbol{\theta}_{exp}\big{\}}^{1:N}\in\mathbb{R}^{N\times 200}$. At each of the denoising timesteps t, FaceTalk predicts the denoised sample and noises it back to timestamp ${t-1}$ :
@@
These smoothed expression deformations $\big{\{}\boldsymbol{\delta}_{exp}^{s}\big{\}}^{1:N}$, along with identity $\boldsymbol{\theta_{id}}$ and predicted expression codes $\big{\{}\boldsymbol{\hat{\theta}}_{exp}\big{\}}^{1:N}$ are passed to the NPHM identity MLP $\big{\{}I\big{\}}$ to obtain smoorhed SDF, from which meshes are extracted using MC [46].
@@ 5 Dataset Creation
where .$\|.\|_{1}$, .$\|.\|_{2}$ and .$\|.\|_{\epsilon}$ refers to L1, L2 and Huber loss respectively, and $n$ refers to number of frames simultaneously optimized. More details regarding aligning the coordinate system of NeRSemble dataset [41] with NPHM can be found in the supplemental document.
Once optimized, next to generate mesh sequences, we uniformly sample points from a 3D grid $\mathcal{P}\in\mathbb{R}^{|X|\times|Y|\times|Z|}$ such that $\mathcal{P}_{xyz}=[p_{x},p_{y},p_{z}]$ is a point in 3d space. These points, along with the identity $\boldsymbol{\theta_{id}}$ and expression codes $\big{\{}\boldsymbol{\theta}^{i}_{\boldsymbol{exp}}\big{\}}_{i=1}^{N}$ are passed to the pretrained NPHM model outputting an SDF; meshes are then extracted with  [46]. In total, our FaceTalk dataset consists of 1000 sequences, an order of magnitude larger than the existing datasets [15].
@@ 6 Results
Unlike template-based methods that rely on LVE (Lip Vertex Error), our method produces SDFs for each frame; extracted mesh topologies thus vary across frames, making LVE inapplicable for evaluation. Thus, we employ LSE-D (Lip Sync Error Distance) [54] for quantitative evaluation of lip synchronization. We evaluate all baselines using this metric. For diversity, we compute FID and KID, as well as diversity scores [56]. Finally, for perceptual evaluation, we perform a user study with 40 participants on a diverse set of 15 unseen audio clips.
Most baseline methods are trained using Vocaset [15]. Thus, for a fair comparison, we construct a hybrid audio test set of 100 sequences, with 25 test audio clips each from Vocaset ($\sim$ 2-3 seconds) and ours ($\sim$ 2-3 seconds), and 50 ($\sim$ 5-7 seconds) from LJSpeech [35]. Test audios were selected from identities not seen during training of any method.
LSE-D metric is then evaluated on this hybrid test set. The FaceTalk dataset consists of 1000 sequences, of which we use 800/100/100 for train/val/test.
@@
For dataset creation, we optimize NPHM expressions for 500 iterations in a group of $n=10$ frames with a step size of 0.001 for iterations $\leq$ 300, and 0.0001 otherwise. We use $\lambda_{sdf}=10$, $\lambda_{temp}=0.1$ and $\lambda_{exp}=0.0025$ for Eq. 16 during optimization. To train our diffusion model, we randomly clip sequences to 2 seconds (48 frames) for efficient minibatching. We train with Adam optimizer with a learning rate of 0.0001, batch size of 32. For diffusion, we use 1000 noising/denoising timestamps, a cosine schedule to add noise to input sequences, and train for 50K iterations. We apply data augmentation to the expression codes by uniformly modulating them.
@@
This is the first work to perform speech-conditioned synthesis for volumetric head motion sequences. We compare with state-of-the-art template-based methods in Fig 4 and Tab. 1. Specifically, we compare against speech-driven animation methods [15, 57, 25, 80], personalized methods  [71], as well as recent emotion-based methods [49, 16]. Our method consistently achieves better lip-audio synchronization while also representing fine-scale facial details like creasing in the nasolabial folds with wider mouth motions. This is confirmed by our perceptual user study in Fig 6.
Method LSE-D $\downarrow$ FID $\downarrow$ KID $\downarrow$
VOCA [15] 13.6191 239.043 0.280
MeshTalk [57] 12.9607 220.172 0.254
FaceFormer [25] 11.9848 215.274 0.222
CodeTalker [80] 11.8054 208.064 0.207
Imitator [71] 11.6119 208.479 0.212
EmoTalk [49] 11.7485 201.311 0.201
EMOTE [16] 11.7192 227.924 0.247
Ours 11.2737 40.692 0.009
@@
Our Method LSE-D $\downarrow$ Diversity $\uparrow$ FID $\downarrow$ KID $\downarrow$
w/o expr. aug. 11.6229 1.61e-8 42.850 0.0092
w/o facial smoothing 11.3488 0.34 43.791 0.0092
w/o FiLM layer 12.029 0.006 47.168 0.0095
w/o expr.-audio align. 13.720 5.17e-8 139.167 0.108
w/o diffusion 11.3217 0.000 41.064 0.0093
Full (Ours) 11.2737 0.34 40.692 0.0092
Table 2: Ablation over model design. Expression augmentation improves diversity, while facial smoothing alleviates inter-frame jitter and improves lip syncronization. Using FiLM conditioning achieves accurate mouth pose and increased diversity. Without expression-audio alignment, the model ignores the audio signal. Without diffusion, the model fails to generate diverse results. Our full model with all components achieves the best results.

### 17 Mar 2024 v2

@@ Abstract
We introduce FaceTalk [^1] Project Page: https://shivangi-aneja.github.io/projects/facetalk, a novel generative approach designed for synthesizing high-fidelity 3D motion sequences of talking human heads from input audio signal. To capture the expressive, detailed nature of human heads, including hair, ears, and finer-scale eye movements, we propose to couple speech signal with the latent space of neural parametric head models to create high-fidelity, temporally coherent motion sequences. We propose a new latent diffusion model for this task, operating in the expression space of neural parametric head models, to synthesize audio-driven realistic head sequences. In the absence of a dataset with corresponding NPHM expressions to audio, we optimize for these correspondences to produce a dataset of temporally-optimized NPHM expressions fit to audio-video recordings of people talking. To the best of our knowledge, this is the first work to propose a generative approach for realistic and high-quality motion synthesis of volumetric human heads, representing a significant advancement in the field of audio-driven 3D animation. Notably, our approach stands out in its ability to generate plausible motion sequences that can produce high-fidelity head animation coupled with the NPHM shape space. Our experimental results substantiate the effectiveness of FaceTalk, consistently achieving superior and visually natural motion, encompassing diverse facial expressions and styles, outperforming existing methods by 75% in perceptual user study evaluation.
Figure 1: Given an input speech signal, we propose a diffusion-based approach to synthesize high-quality and temporally consistent 3D motion sequences of high-fidelity human heads as neural parametric head models. Our method can generate a diverse set of expressions (including wrinkles and eye blinks) and the generated mouth motion is temporally synchronized with the given audio signal.
@@ 1 Introduction
Modeling 3D animation of humans has a wide range of applications in the realm of digital media, including animated movies, computer games, and virtual assistants. In recent years, there have been numerous works proposing generative approaches for motion synthesis of human bodies, enabling the animation of human skeletons conditioned on various signals such as action [29, 53, 3], language [1, 54, 72, 85, 41, 4] and music [75, 2]. While human faces are critical to synthesis of humans, generative synthesis of 3D faces in motion has focused on 3D morphable models (3DMMs) leveraging linear blendshapes [7, 45] to represent head motion and expression. Such models characterize a disentangled space of head shape and motion, but lack the capacity to comprehensively represent the complexity and fine-grained details of human face geometry in motion (e.g., hair, skin furrowing during motion, etc.).
Thus, we propose to represent animated head sequences with a volumetric 3D head representation, leveraging the expressive representation space of neural parametric head models (NPHMs) [27, 28]. NPHMs offer a flexible representation capable of handling complex and irregular facial expressions (e.g., blinking, skin creasing), along with a high-fidelity shape space including the head, hair, and ears, making them a much more suitable choice for face animation. We address the challenging task of creating an audio-conditional generative animation model for this volumetric representation.
@@ 3 Preliminaries
In contrast to traditional 3D morphable models, which remain limited in expressive detail, we employ the Neural Parametric Head Model (NPHM) representation [27, 28]. Similar to traditional 3DMMs, NPHM disentangles a human face into an identity and expression space; however, the shape and expression space are volumetric, enabling expressive capture of details such as fine-scale eye movements, hair, ears, etc. NPHM uses two auto-decoder style neural networks (a) Identity Network ${}\big{\{}\mathcal{I}\big{\}}$ to represent the overall facial shape and (b) Expression Network ${}\mathcal{E}\big{\}}$ to represent the facial movements such as jaw pose, wrinkles, eyeblinks, etc., jointly denoted as $\mathcal{F}=\big{\{}\mathcal{I},\mathcal{E}\big{\}}$. The identity and expressions latent codes $\big{\{}\boldsymbol{{\theta}_{id}},\boldsymbol{{\theta}_{exp}}\big{\}}$ define the facial shape and expression respectively. Unlike 3DMMs which predict a fixed-topology mesh that can not model fine surface details, NPHM can handle different hairstyles, wrinkles and complex facial expressions. NPHM represents identities with a signed distance field (SDF) in canonical space and models the expressions as deformations onto the facial region. Mathematically, it can be defined as:
@@
Diffusion models are a class of generative models that consist of a forward and reverse process. The forward process converts the original structured data distribution into Gaussian noise, modeled following a fixed Markov chain as:
@@ 4 Method
FaceTalk performs high-fidelity and temporally consistent generative synthesis of motion sequences of heads, conditioned on audio signal. In order to characterize complex face motions and fine-scale movements, we synthesize realistic heads in the latent expression space of a parametric model for volumetric head representations, i.e., neural parametric head models (NPHMs) [27]. We, thus, develop a speech-conditioned latent diffusion model to synthesize temporally coherent head expression sequences that are coupled with the NPHM shape space to produce complex, realistic head animations of different identities. An overview of our approach is illustrated in Fig. 2.
@@
We train our expression decoder network on optimized NPHM expression sequences $\big{\{}\boldsymbol{\theta}_{exp}\big{\}}^{1:N}$ (obtained in Section 5), where $N$ refers to number of frames in a sequence. We train a diffusion-based stacked multi-layer transformer [76] decoder network to synthesize facial expressions in the latent space of the NPHM model. During training, following forward diffusion (Eq. 3) we add noise for a randomly sampled diffusion timestamp $t\sim\mathtt{Uniform}(0,T)$ to create noisy expression codes $\big{\{}\boldsymbol{\theta}_{exp}\big{\}}^{1:N}_{t}$. These noisy expression codes are then projected to the latent space of our model via a linear layer, followed by a stack of transformer decoder blocks, and then projected back to the original NPHM space using another linear layer. To embed diffusion timestamp in the latent space of our model, we apply sinusoidal embedding and process it through a three-layer MLP. Next, to fuse diffusion timestamp into the model, we use a one-layer FiLM (feature-wise linear modulation) network [52] between the multi-head self-attention, multi-head cross-attention and feedforward layers of the transformer decoder block, which is critical for the model to produce high-quality expressions as we show in the results. We leverage a look-ahead binary target mask $\mathcal{T}\in\mathbb{R}^{N\times N}$ in the multi-head self-attention layers to prevent the model from peeking into the future expression codes. Mathematically, it can be written as:
@@
where $\mathcal{T}_{ij}$ refers to the $(i,j)^{th}$ element of the matrix and $1\leq i,j\leq N$. The audio features $\boldsymbol{A}^{1:N}$ are fused into the network via the multi-head cross-attention layers. We leverage the expression-audio alignment mask $\mathcal{M}$ to fuse the audio features into the network to ensure that mouth poses learned by the expression codes are consistent with the speech signal. The binary mask $\mathcal{M}\in\mathbb{R}^{N\times N}$ is Kronecker delta function $\delta_{ij}$ such that the audio features for $i^{th}$ timestamp attend to expression features at the $j^{th}$ timestamp if and only if $i=j$. Mathematically,
@@
Unlike other domains (like text-to-motion) where the synthesized motion can vary dramatically, speech-driven 3D facial animation is more constrained, requiring precise mouth alignment with the audio signal, making it prone to overfitting. To alleviate this, we propose a novel augmentation strategy capable of producing diverse expressions for a given audio signal. Our key insight is to augment the dataset to generate different expression codes for the same speech signal by randomly amplifying and suppressing its magnitude. Specifically, we randomly sample modulation factor $r$ within the bounds $[a,b]$ as:
$r\sim\mathtt{Uniform}(a,b),$ (8)
and scale the expression codes as $\big{\{}r\times\boldsymbol{\theta}_{exp}\big{\}}^{1:N}$. We show in the results this augmentation strategy helps the model synthesize diverse expressions.
@@
FaceTalk uses a diffusion-based framework to learn to synthesize NPHM expression sequences of $N$ frames $\big{\{}\boldsymbol{\theta}_{exp}\big{\}}^{1:N}\in\mathbb{R}^{N\times 200}$. At each of the denoising timestep t, FaceTalk predicts the denoised sample and noises it back to timestamp ${t-1}$ :
@@
These smoothed expression deformations $\big{\{}\boldsymbol{\delta}_{exp}^{s}\big{\}}^{1:N}$, along with identity $\boldsymbol{\theta_{id}}$ and predicted expression codes $\big{\{}\boldsymbol{\hat{\theta}}_{exp}\big{\}}^{1:N}$ are passed to the NPHM identity MLP $\big{\{}I\big{\}}$ to obtain smoothed SDF, from which meshes are extracted using MC [47].
@@ 5 Dataset Creation
where .$\|.\|_{1}$, .$\|.\|_{2}$ and .$\|.\|_{\epsilon}$ refers to L1, L2 and Huber loss respectively, and $n$ refers to number of frames simultaneously optimized. More details regarding aligning the coordinate system of NeRSemble dataset [42] with NPHM can be found in the supplemental document. Once optimized, next to generate mesh sequences, we uniformly sample points from a 3D grid $\mathcal{P}\in\mathbb{R}^{|X|\times|Y|\times|Z|}$ such that $\mathcal{P}_{xyz}=[p_{x},p_{y},p_{z}]$ is a point in 3d space. These points, along with the identity $\boldsymbol{\theta_{id}}$ and expression codes $\big{\{}\boldsymbol{\theta}^{i}_{\boldsymbol{exp}}\big{\}}_{i=1}^{N}$ are passed to the pretrained NPHM model outputting an SDF; meshes are then extracted with  [47]. In total, our FaceTalk dataset consists of 1000 sequences, an order of magnitude larger than the existing datasets [15].
@@ 6 Results
Unlike template-based methods that rely on LVE (Lip Vertex Error), our method produces SDFs for each frame; extracted mesh topologies thus vary across frames, making LVE inapplicable for evaluation. Thus, we employ LSE-D (Lip Sync Error Distance) [56] for quantitative evaluation of lip synchronization. For diversity, we compute FID and KID, as well as diversity scores [58]. As FID/KID may not reflect all quality considerations, when samples are limited, we include two established quality metrics: (1) FIQA (Face Image Quality Assessment) [49] (2) VQA (Video Quality Assessment) [82]. FIQA quantifies image quality for face recognition and similarity to in-the-wild real faces and VQA measures overall video quality in terms of distortions/semantics/aesthetics. Finally, for perceptual evaluation, we perform a user study with 40 participants on a diverse set of 15 unseen audio clips. Most baseline methods are trained using Vocaset [15]. Thus, for a fair comparison, we construct a hybrid audio test set of 100 sequences, with 25 test audio clips each from Vocaset ($\sim$ 2-3 seconds) and ours ($\sim$ 2-3 seconds), and 50 ($\sim$ 5-7 seconds) from LJSpeech [36]. Test audios were selected from identities not seen during training of any method. The metrics are reported on this hybrid test set.
@@
For dataset creation, we optimize NPHM expressions for 500 iterations in a group of $n=10$ frames with a step size of 0.001 for iterations $\leq$ 300, and 0.0001 otherwise. We use $\lambda_{sdf}=10$, $\lambda_{temp}=0.1$ and $\lambda_{reg}=0.0025$ for Eq. 16 during optimization. To train our diffusion model, we randomly clip sequences to 2 seconds (48 frames) for efficient minibatching. We train with Adam optimizer with a learning rate of 0.0001. For diffusion, we use 1000 noising/denoising timestamps, a cosine schedule to add noise to input sequences. We apply data augmentation to the expression codes by uniformly modulating them.
@@
This is the first work to perform speech-conditioned synthesis for volumetric head motion sequences. We compare with state-of-the-art template-based methods in Fig 4 and Tab. 1. Specifically, we compare against speech-driven animation methods [15, 59, 25, 83], personalized methods  [73], as well as recent emotion-based methods [51, 16]. Our approach more accurately captures subtle mouth movements compared to less expressive baselines, resulting in better FID/KID. These scores were evaluated only on mouth region crops to avoid bias towards rest of facial geometry. Our method consistently achieves better lip-audio synchronization while also representing fine-scale facial details like creasing in the nasolabial folds with wider mouth motions. This is confirmed by our perceptual user study in Fig 6.
Method LSE-D $\downarrow$ FID $\downarrow$ KID $\downarrow$ FIQA $\uparrow$ VQA $\uparrow$
VOCA [15] 13.6191 239.043 0.280 33.36 0.4251
MeshTalk [59] 12.9607 220.172 0.254 36.14 0.4855
FaceFormer [25] 11.9848 215.274 0.222 38.24 0.5227
CodeTalker [83] 11.8054 208.064 0.207 38.38 0.5274
Imitator [73] 11.6119 208.479 0.212 38.44 0.5384
EmoTalk [51] 11.7485 201.311 0.201 31.80 0.5037
EMOTE [16] 11.7192 227.924 0.247 39.44 0.5149
Ours 11.2737 40.692 0.009 45.75 0.6145
@@
Our Method LSE-D $\downarrow$ Diversity $\uparrow$
w/o expr. aug. 11.6229 1.61e-8
w/o facial smoothing 11.3488 0.34
w/o FiLM layer 12.029 0.006
w/o expr.-audio align. 13.720 5.17e-8
w/o diffusion 11.3217 0.000
Full (Ours) 11.2737 0.34
Table 2: Ablation over model design. Expression augmentation improves diversity, while facial smoothing alleviates inter-frame jitter. Using FiLM conditioning achieves accurate mouth pose and increased diversity. Without expression-audio alignment, the model ignores the audio signal. Without diffusion, the model fails to generate diverse results. Our full model with all components achieves the best results.